Opened 5 years ago

Closed 4 years ago

#519 closed flight processing (fixed)

GB12/04, flight day 083/2014, Eaves Wood

Reported by: emca Owned by: knpa
Priority: immediate Milestone:
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by tec)

Data location: /users/rsg/arsf/arsf_data/2014/flight_data/uk/GB12_04-2014_083_Eaves_Wood

Data arrived from ARSF via network transfer on 26/03/2014

GB12/04 Scientific objective: developing the use of multi-temporal LiDAR data for quantifying forest dynamics

GB12/04 Priority: a4m

PI: Alan Blackburn

Notes: Full waveform data present

Sensors:

Fenix (Requested, Submitted to NEODC @ 20/04/2015)
LiDAR (Requested, Submitted to NEODC @ 20/04/2015)
FW LiDAR (Requested, Submitted to NEODC @ 20/04/2015)

Change History (23)

comment:1 Changed 5 years ago by emca

Raw full waveform data for 5 out of 24 flightlines. Checking with Ops what happened to the other lines.

comment:2 Changed 5 years ago by knpa

Basestation position through PPP is:

LAT 54 04 33.18161
LONG -2 18 31.77521
HEIGHT 284.212

comment:3 Changed 5 years ago by knpa

Navigation processing complete.

comment:4 Changed 5 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

comment:5 Changed 5 years ago by knpa

Doing FENIX processing

There are no dark frames for this flight.

comment:6 Changed 5 years ago by knpa

Files 14, 20 and 23 didn't have properly overlapping nav files. Forced it to use those navs with: nav2_args_extra = -nav

comment:7 Changed 5 years ago by knpa

Found SCTs. They seem to be mostly ~0.95 except when they're not...
Not sure what happened with 11, it's a line where Fenix crashed.

FENIX
-1: 0.95
-2: 0.94
-3: -0.06
-4: 0.94
-5: 0.95
-6: 0.89
-7: 0.94
-8: 0.96
-9: 0.95
-10: 0.95
-11: -24.02
-12: 0.93
-13: -0.04
-14: 0.96
-15: 0.95
-16: 0.96
-17: 0.96
-18: 0.94
-19: 0.99
-20: 0.96
-21: 0.95
-22: 0.94
-23: 0.97
-24: 0.96
-25: 0.93

comment:8 Changed 5 years ago by knpa

Dark frames issue ongoing - cannot proceed any further with hyperspectral for now.

comment:9 Changed 5 years ago by knpa

Going to use dark frames from 152a.
Has same integration time, fps, binning.

Can salvage VNIR but not SWIR.

comment:10 Changed 5 years ago by dac

Started LiDAR processing

comment:11 Changed 5 years ago by dac

Finished LiDAR processing and waiting for delivery check.

Single roll/pitch offset used for all lines:
Roll: -0.0037
Pitch: -0.00092

Processed available waveform data (5 lines) - waiting to hear back from Ops on waveform data for other flightlines.

DEM location checked in GoogleEarth.

comment:12 Changed 5 years ago by lah

Started lidar delivery check.

Only 5 flightlines of full waveform data present (see previous comments).

Full waveform.pdf missing from doc folder and no fw extractions.
.sup file is missing from navigation folder.
Should .tif file be present in screenshots folder? No vectors screenshot.

Typos in Read_Me.pdf: 'incase' pg4 ln 5, 'are' pg4 ln 16.

Text files are ascii, but not clrf terminated.

Too many points are classified as noise, particularly on steep gradients and within trees.

comment:13 Changed 5 years ago by dac

Have corrected most of the problems:

  • Copied full_waveform_lidar.pdf from /users/rsg/arsf/doc/data_quality_reports/full_waveform_lidar.pdf
  • Checked with mark1, .sup file no longer required (have updated wiki).
  • Removed tif file from screenshots.
  • Have changes incase to two words.
  • Have changed the following sentence: "The intensity has values between 0 and 255 and are affected by an automatic gain control so may not be suitable for using as scientific data." to "The intensity has values between 0 and 255 and is affected by an automatic gain control so may not be suitable for using as scientific data." Also needs changing on the readme template.
  • Converted ASCII data to use CRLF line terminators.

Still waiting to hear back on full waveform data for remaining strips - will look into.

Flight lines cover quarry with steep sides, some points on these have been erroneously classified as noise - looking into.

comment:14 Changed 5 years ago by dac

Have revisited classification. Given the anticipated time required to manually correct points classified as noise around quarry and scientific objectives of quantifying forest dynamics have left classification as is.

A number of LiDAR lines were flown over the same area, only 5 of the lines were flown with waveform data. All expected LiDAR data are present.

Last edited 5 years ago by dac (previous) (diff)

comment:15 Changed 5 years ago by knpa

Created HS delivery.

Currently removing SWIR portion using /users/rsg/arsf/arsf_data/2014/misc/2014_cal_feb_gloucester/software/fenix_splitter/splitfile

comment:16 Changed 5 years ago by dac

Finished and sent LiDAR data on 27/08/2014 to:

Alan Blackburn
Lancaster Environment Centre,
Lancaster University,
Lancaster,
LA1 4YQ

comment:17 Changed 5 years ago by knpa

There were problems with SWIR and VNIR splitting.

Mark1 is currently looking into fixing the scripts, so waiting on that.

comment:18 Changed 5 years ago by dac

Fenix Processing
Took over from knpa.
Fixed VNIR/SWIR splitting program so map info is copied across correctly. Split mapped files and removed SWIR bands (level1b data and masks were already split).
Ran through delivery check, fixing problems as required.
Ready to deliver pending second check.

comment:19 Changed 5 years ago by dap

Beginning Fenix Delivery Check

comment:20 Changed 5 years ago by dap

Fenix Delivery Check complete.

The following issues were found with the delivery:

  • "contact us using the details on page 12" - no contact details on this page. Fixed this by inserting dynamic page reference to the contact details section.
  • Broken link in read me - "Japan's Ground Data System" - fixed read me and template
  • Broken link - "ASTER GDEM Validation Summary Report (pdf)" - fixed read me and template
  • Broken link - "https://lpdaac.usgs.gov/lpdaac/products/aster_policies" - fixed read me and template
  • Broken link - NEODC website - fixed in read me and template
  • Removed flightlines/mapped/f08311a3b_mapped_osng.bil.aux.xml and flightlines/mapped/f08311b3b_mapped_osng.bil.aux.xml.
  • logsheet was incorrectly named "GB12_04-2014_83-Eaves_Wood.pdf", have renamed to "GB12_04-2014_083-Eaves_Wood.pdf".
  • proj_tidy.sh pointed out f08311[a-b]1b.* as unrecognised files, ignoring this as was mentioned in read me.

Dataset at GB12_04-083-hyperspectral-20140721 will be ready for delivery once zip_mapped has completed.

Last edited 5 years ago by dap (previous) (diff)

comment:21 Changed 5 years ago by dac

Fenix Delivery
Finished and sent Fenix data on 22/10/2014 to:

Alan Blackburn
Lancaster Environment Centre,
Lancaster University,
Lancaster,
LA1 4YQ

comment:22 Changed 4 years ago by tec

Archiving
Started

comment:23 Changed 4 years ago by tec

  • Description modified (diff)
  • Resolution set to fixed
  • Status changed from assigned to closed

Archiving
Removed LiDAR processing/als50/las_alsproc as it can be regenerated
Removed Hyperspectrial processing/hyperspectrial/flightlines as it can be regenerated

Finished

Note: See TracTickets for help on using tickets.